OpenStudy (anonymous):

clear fractions or decimals, solve, and check 2/3+1/4t=2; show work

OpenStudy (thomas5267):

Can you use TeX to represent your question as there are many interpretations of "2/3+1/4t=2"

OpenStudy (saifoo.khan):

\[\frac 23 + \frac14 t = 2\] Like that right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Assuming the way saifoo.khan has structured it, you can fraction bust by multiplying everything by 12 to get: 8 + 3t = 24 then 3t = 16 and t = 16/3 Check: (2/3) + (1/4)(16/3) = 2 (2/3) + (16/12) = 2 (8/12) + (16/12) = 2 24/12 = 2 2=2

OpenStudy (saifoo.khan):

\[\frac14t = 2 - \frac23\]\[\frac14 t = \frac{2(3) - 2}{3}\]\[\frac14t = \frac{4}{3}\]\[t = 4\times \frac{4}{3}\]